OpenStudy (anonymous):

Algebra 1 help. Will give medal and fan!:)

OpenStudy (anonymous):

4. Esmeralda is graphing a polynomial function as a parabola. Before she begins graphing it, explain how to find the vertex. Make sure you include how to determine if it will be a maximum or minimum point. Use an example quadratic function to help you explain and provide its graph.

OpenStudy (anonymous):

@hl147 please help?

OpenStudy (anonymous):

Im sorry im not sure how to do this question

OpenStudy (anonymous):

Do you know anyone who does? Please it's my last one

OpenStudy (anonymous):

try asking people who have a higher number and a lot of fans

OpenStudy (anonymous):

i also suggest searching your question on google or youtube

OpenStudy (anonymous):

I don't have time to look for people like that. I just started today and I need to turn this in in an hour. I have tried to figure it out on my own and icant.

OpenStudy (anonymous):

oh okay well try this website http://mathbitsnotebook.com/Algebra1/Quadratics/QDVertexForm.html

OpenStudy (anonymous):

i think you use the vertex form to find the vertex

OpenStudy (anonymous):

OK well thanks. I'll try. You still get the medal and fan.

OpenStudy (anonymous):

aww thanks you didn't have to , and sorry i learned this a while away ago

OpenStudy (anonymous):

Nah its fine lol

OpenStudy (anonymous):

Graphing a Quadratic Function in Vertex Form: 1. Start with the function in vertex form: y = a(x - h)2 + k y = 3(x - 2)2 - 4 2. Pull out the values for h and k. If necessary, rewrite the function so you can clearly see the h and k values. (h, k) is the vertex of the parabola. Plot the vertex. y = 3(x - 2)2 + (-4) h = 2; k = -4 Vertex: (2, -4) 3. The line x = h is the axis of symmetry. Draw the axis of symmetry. x = 2 is the axis of symmetry

OpenStudy (anonymous):

this is an example

OpenStudy (anonymous):

Ok. Yeah i got it.

OpenStudy (anonymous):

ok good